translate and solve eight less than three times a number is equal to the sum of the number and four

Answer by ewatrrr(24785)

 
Hi,
translate and solve eight less than three times a number is equal to the sum of the number and four
Let x represent the number
Question states***
3x-8 = x+ 4
Solving for x
2x = 12
x = 6
CHECKING our Answer***
18-8 = 6 + 4